# Unveil Multi-Picture Descriptions for Multilingual Mild Cognitive Impairment Detection via Contrastive Learning

### v2

@@ I Introduction
Traditional cognitive assessments rely on blood tests, neuroimaging, and in-clinic tests like the Mini-Mental State Examination (MMSE). While effective, these methods are time-consuming and labor-intensive. In contrast, AI-driven speech processing has emerged as a tool for early MCI detection through non-invasive and cost-effective screening. These processing systems can automatically identify biomarkers related to cognitive decline, such as hesitation gaps, reduced fluency, and changes in word richness [4].
@@ II-A Multilingual Studies in Detecting Cognitive Decline
Speech and language processing for cognitive decline detection spans several decades, with most methods focusing on English speakers. Two widely used speech data are the ADReSS challenge and Pitt Corpus [5]. Speech processing combines acoustic and linguistic features with machine learning classifiers. Studies have reported accuracy ranging from 72.9% to 89.6% depending on models [14, 15]. A few studies include more languages. The ADReSS-M challenge includes English and Greek speakers for dementia detection, achieving accuracy up to 87% [16]. Pérez-Toro et al. have included English, German, and Spanish speakers for Alzheimer’s detection [17] and achieved UAR of 73%. Furthermore, the TAUKDIAL-2024 challenge focuses on MCI detection, and the established baseline achieves 59.18% UAR. Other reported UARs are between 49.4% and 56.18% [8, 9]. These results imply the challenges in developing highly performing and generalizable models for multilingual MCI detection.
@@ II-B Overfitting and Spurious Correlation
Recent work by Liu et al. [11] raised important questions about the reliability of reported high accuracy in dementia detection. Their analysis of the Pitt Corpus revealed that models could achieve good results by relying solely on silence segments, which is the Clever Hans effect. This finding suggests that models might be exploiting dataset-specific shortcuts, leading to spurious correlations rather than learning true causal patterns. Spurious correlations can come from imbalanced and noisy datasets. The consequence of spurious correlations is bias against minorities and over-reliance on data shortcuts, limiting real-world application deployment. While rectifying the weakness of relying on spurious correlation is crucial for clinical applications, it remains a challenge in adapting existing algorithms in medical routines. To address the challenges, we adopt PoE to mitigate reliance on spurious features and reduce performance disparities between subgroups.
@@ V-A Preprocessing
For preprocessing speech inputs, we first removed the noisy background and applied a 16 kHz sampling rate to the audio with a single (mono) channel. Speech was transcribed using the Whisper large-v3111https://huggingface.co/openai/whisper-large-v3 automatic speech recognition (ASR) model to convert audio recordings into text. We maintained the default feature dimension settings for pre-trained models.
@@ V-B Feature Extraction
Speech Encoder: We process each speech input $S_{i}$ through a speech encoder $f_{S}$ to obtain speech embeddings $f_{S}(S_{i})$. Specifically, we use the pre-trained Whisper-tiny model [^2] https://huggingface.co/openai/whisper-tiny [19], which extracts features from a transformer-based architecture.
Acoustic Features: The handcrafted acoustic features supplement previous speech encoder features. We extract a comprehensive set of acoustic features $f_{A}(S_{i})$ using the DisVoice [^3] https://github.com/jcvasquezc/DisVoice package [20]. These handcrafted features are phonation characteristics, phonological information, articulation dynamics, and prosodic elements, which are relevant to identifying the MCI speech patterns.
Text Encoder: For each transcription $T_{i}$, linguistic embeddings $f_{T}(T_{i})$ are obtained using the pre-trained language model. We use the BERT-base [^4] https://huggingface.co/google-bert/bert-base-uncased [21] as the text encoder.
@@ V-D Product of Experts for Multimodal Concatenation
where [;][;][ ; ] denotes concatenations. $\mathrm{FFN}_{M}$ is a feed-forward network to transform multimodal concatenation. $z_{M}$ denotes the predicted class logits from the multimodal concatenation.
@@
Next, we aim to combine these predictions efficiently, where the PoE operates in the log space. Rather than performing element-wise multiplication of probabilities, we sum their logarithms values (element-wise addition) because log-space operation is more efficient than probability multiplication.
@@ VI-B Implementation Details
To ensure robust evaluation given small and imbalanced data, we trained our framework with stratified k-fold cross-validation (k=10). Hyperparameters include a learning rate of 1e-5, a batch size of 16, L2 regularization penalty$=0.01$, and the Adam optimizer. The framework is trained for 10 epochs and implemented in PyTorch on an A100 GPU.
@@ VI-C Evaluation Metrics
[…]
@@ VII-A Baseline Performance
TABLE II: Evaluation across each unimodality with CL and IE. Results include entire data and language and gender subgroup performances. Green highlights values above the baseline in each column.
@@ VII-B After Adding CL and IE to Unimodality
CL and IE are more effective for the text modality than for the speech modality. As shown in Table II, adding CL to $z_{T}$ increases UAR from 68.1% baseline to 71.8% (+3.7%). The addition of IE further improves it by 5.0% (73.0%). The F1 score improves from 80.6% to 82.0% after adding both CL and IE. In contrast, speech models show no substantial improvements. These indicate that CL can effectively enhance the model’s ability to discriminate linguistic features and semantic differences from multi-picture descriptions. The inclusion of IE provides more picture context in addition to text.
@@ VII-C After Combining All Modalities
CE and IE integration with PoE improves multimodal baseline performance. Shown in Table III, our multimodal baseline $Z_{M}$ achieves UAR of 72.4%, outperforming all unimodal approaches. Incorporating PoE further improves UAR to 73.6%, since PoE encourages individual modalities to contribute collectively to make the final decision. This process aims to enhance model generalization. When we combine PoE, CL and IE, we observe additional improvements compared with $Z_{M}$: increases in UAR by +2.8% (72.4% to 75.2%) and F1 score by +2.2% (81.3% to 83.5%). These results imply that the combination of CL and IE with PoE leads to an improved framework for MCI detection.
TABLE III: Comparison of multimodal baseline ($z_{M}$), $z_{M}$ with PoE, and their combinations with CL and IE. Bold values are where $z_{M}$ + PoE outperforms $z_{M}$. Green highlights indicate improvements above $z_{M}$ + PoE.
@@
PoE reduces disparity across subgroups. The inclusion of PoE leads to improvements either in UAR or F1 score compared to the underperforming baseline subgroups. In the gender subgroup, female has UAR increase of +3.2% (70.0% to 73.2%) and F1 score increase of +3.7% (78.4% to 82.1%). The male performance significantly reduces, thereby reducing the gender disparity. Similarly, since Chinese UAR reduces while English remains stable, narrowing the gap between the previously stronger Chinese baseline. These results exhibit that PoE can effectively elevate worse-performing subgroups while maintaining outcomes from better-performing subgroups.
@@ VII-D The Impact of Contrastive Learning on Other Models
To further understand the effectiveness of adding CL, we evaluate it across multiple linguistic and speech feature extraction models. Linguistic models include BERT, RoBERTa [^5] https://huggingface.co/FacebookAI/roberta-base [23], and XLM [^6] https://huggingface.co/FacebookAI/xlm-roberta-base [24]. Speech models include Whisper, HuBERT [^7] https://huggingface.co/facebook/hubert-base-ls960 [25], and Wav2Vec2888https://huggingface.co/facebook/wav2vec2-base [26].
Figure 2: Performance impact of contrastive learning on different models. A-B UAR and F1 scores for linguistic models (BERT, RoBERTa, XLM) with and without CL. C-D UAR and F1 scores for speech models (Whisper, HuBERT, Wav2Vec2) with and without CL.
@@ VIII Limitation
Our work primarily uses the supervised contrastive learning loss and PoE loss functions. We hypothesize that the results will generalize beyond the loss functions we discussed. Future validation is needed to verify this. Other questions remain open: we explore a limited set of MCI-relevant handcrafted features. Future work will analyze more handcrafted features and how they are correlated with the performance of our framework. We will also compare with other multimodal fusion strategies to compare their performance with ours. In addition, the small and imbalanced data limits the highly performing features and models, which requires future data collection and benchmark efforts.

### v5

@@ I Introduction
Traditional cognitive assessments rely on blood tests, neuroimaging, and in-clinic tests like the Mini-Mental State Examination (MMSE). While effective, these methods are time-consuming and labor-intensive. In contrast, AI-driven speech processing has emerged as a tool for early MCI detection through non-invasive and cost-effective screening. These processing systems can automatically identify biomarkers related to cognitive decline, such as hesitation gaps, reduced fluency, vocal tremors, and changes in word richness [4, 5].
@@ II-A Multilingual Studies in Detecting Cognitive Decline
Speech and language processing for cognitive decline detection spans several decades, with most methods focusing on English speakers. Two widely used datasets are the ADReSS challenge and Pitt Corpus [6]. Speech processing combines acoustic and linguistic features with machine learning classifiers. Studies have reported accuracy ranging from 72.9% to 89.6% depending on models [15, 16]. A few studies include more languages. ADReSS-M challenge includes English and Greek speakers for dementia detection, achieving accuracy up to 87% [17]. Pérez-Toro et al. have included English, German, and Spanish speakers for Alzheimer’s detection [18] and achieved UAR of 73%. Furthermore, the TAUKDIAL-2024 challenge focuses on MCI detection, and the established baseline achieves 59.18% UAR. Other reported UARs are between 49.4% and 56.18% [9, 10]. These results imply the challenges in developing highly performing and generalizable models for multilingual MCI detection.
@@ II-B Overfitting and Spurious Correlation
Recent work by Liu et al. [12] raised important questions about the reliability of reported high accuracy in dementia detection. Their analysis of the Pitt Corpus revealed that models could achieve good results by relying solely on silence segments, which is the Clever Hans effect. This finding suggests that models might be exploiting dataset-specific shortcuts, leading to spurious correlations rather than learning true causal patterns. Spurious correlations can come from imbalanced and noisy datasets. The consequence of spurious correlations is bias against minorities and over-reliance on data shortcuts, limiting real-world application deployment [19]. While rectifying the weakness of relying on spurious correlation is crucial for clinical applications, it remains a challenge in adapting existing algorithms to medical routines. To address the challenges, we adopt PoE to mitigate reliance on spurious features and reduce performance disparities between subgroups [20, 21].
@@ V-A Preprocessing
For preprocessing speech inputs, we first removed the noisy background and applied a 16 kHz sampling rate to the audio with a single (mono) channel. Speech was transcribed using the Whisper large-v3 [23] automatic speech recognition (ASR) model to convert audio recordings into text. We maintained the default feature dimension settings for pre-trained models.
@@ V-B Feature Extraction
Speech Encoder: We process each speech input $S_{i}$ through a speech encoder $f_{S}$ to obtain speech embeddings $f_{S}(S_{i})$. Specifically, we use the pre-trained Whisper-tiny model [23], which extracts features from a transformer-based architecture.
Acoustic Features: The handcrafted acoustic features supplement previous speech encoder features. We extract a comprehensive set of acoustic features $f_{A}(S_{i})$ using the DisVoice package [24]. These handcrafted features are phonation characteristics, phonological information, articulation dynamics, and prosodic elements, which are relevant to identifying the MCI speech patterns.
Text Encoder: For each transcription $T_{i}$, linguistic embeddings $f_{T}(T_{i})$ are obtained using the pre-trained language model. We use the BERT-base[25] as the text encoder.
@@ V-D Product of Experts for Multimodal Concatenation
where $[;]$ denotes concatenations. $\mathrm{FFN}_{M}$ is a feed-forward network to transform multimodal concatenation. $z_{M}$ denotes the predicted class logits from the multimodal concatenation.
@@
Next, we aim to combine these predictions efficiently, where the PoE operates in the log space. Rather than performing element-wise multiplication of probabilities, we sum their logarithm values (element-wise addition) because log-space operation is more efficient than probability multiplication.
@@ VI-B Implementation Details
To ensure robust evaluation given small and imbalanced data, we trained our framework with stratified k-fold cross-validation (k=10). Hyperparameters include a learning rate of 1e-5, a batch size of 16, L2 regularization penalty=0.01=0.01, and the Adam optimizer. The framework is trained for 10 epochs and implemented in PyTorch on an A100 GPU.
@@ VI-C Evaluation Metrics
We adhere to evaluation metrics required by the challenge for binary classification: UAR and F1 score: $\text{UAR}=\frac{\sigma+\rho}{2}$ and $\text{F1}=\frac{2\pi\rho}{\pi+\rho}$, where $\sigma=\frac{\text{TN}}{\text{TN}+\text{FP}}$ represents specificity, $\rho=\frac{\text{TP}}{\text{TP}+\text{FN}}$ represents sensitivity, and $\pi=\frac{\text{TP}}{\text{TP}+\text{FP}}$ represents precision. TP, TN, FP, and FN denote true positives, true negatives, false positives, and false negatives, respectively.
@@ VII-A Baseline Performance
TABLE II: Results of unimodality with CL and IE, for entire dataset and each language and gender subgroup. Green values indicate performance gain over baseline.
@@ VII-B After Adding CL and IE to Unimodality
CL and IE are more effective for the text modality than for the speech modality. As shown in Table II, adding CL to $z_{T}$ increases UAR from 68.1% baseline to 71.8% (+3.7%). The addition of IE further improves it by 5.0% (73.0%). F1 score improves from 80.6% to 82.0% after adding both CL and IE. In contrast, speech models show no substantial improvements. These indicate that CL can effectively enhance the model’s ability to discriminate linguistic features and semantic differences from multi-picture descriptions. The inclusion of IE provides more picture context in addition to text.
@@ VII-C After Combining All Modalities
CE and IE integration with PoE improves multimodal baseline performance. Shown in Table III, our multimodal baseline $Z_{M}$ achieves UAR of 72.4%, outperforming all unimodal approaches. Incorporating PoE further improves UAR to 73.6%, since PoE encourages individual modalities to contribute collectively to make the final decision. This process aims to enhance model generalization. When we combine PoE, CL, and IE, we observe additional improvements compared with $Z_{M}$: increases in UAR by +2.8% (72.4% to 75.2%) and F1 score by +2.2% (81.3% to 83.5%). These results imply that the combination of CL and IE with PoE leads to an improved framework for MCI detection.
TABLE III: Comparison of multimodal baseline ($z_{M}$), with PoE, CL and IE. Bold values indicate $z_{M}$ + PoE outperforms $z_{M}$. Green highlights indicate improvements over $z_{M}$ + PoE.
@@
PoE reduces disparity across subgroups. The inclusion of PoE leads to improvements either in UAR or F1 score compared to underperforming baseline subgroups. In the gender subgroup, female has UAR increase of +3.2% (70.0% to 73.2%) and F1 score increase of +3.7% (78.4% to 82.1%). The male performance significantly reduces, thereby reducing the gender disparity. Similarly, since Chinese UAR reduces while English remains stable, the gap between the previously stronger Chinese baseline narrows. These results exhibit that PoE effectively elevates worse-performing subgroups while maintaining outcomes from better-performing subgroups.
@@ VII-D The Impact of Contrastive Learning on Other Models
To further understand the effectiveness of adding CL, we evaluate it across multiple linguistic and speech feature extraction models. Linguistic models include BERT, RoBERTa [^1] https://huggingface.co/FacebookAI/roberta-base [27], and XLM [^2] https://huggingface.co/FacebookAI/xlm-roberta-base [28]. Speech models include Whisper, HuBERT [^3] https://huggingface.co/facebook/hubert-base-ls960 [29], and Wav2Vec2444https://huggingface.co/facebook/wav2vec2-base [30].
Figure 2: Performance impact of contrastive learning on different models. A-B: UAR and F1 scores for linguistic models (BERT, RoBERTa, XLM) with and without CL. C-D: UAR and F1 scores for speech models (Whisper, HuBERT, Wav2Vec2) with and without CL.
@@ VIII LIMITATION
Our work primarily uses the supervised contrastive learning loss and PoE loss functions. We hypothesize that the results will generalize beyond the loss functions we discussed. Future validation is needed to verify this. Other questions remain open: we explore a limited set of MCI-relevant handcrafted features. Future work will analyze more handcrafted features and how they are correlated with the performance of our framework. We will also compare with other multimodal fusion strategies to compare their performance with ours. In addition, the small and imbalanced data limit the highly performing features and models, which require future data collection and benchmark efforts.
@@ IX Conclusion
X Acknowledgement
This research is funded by the US National Institutes of Health National Institute on Aging R01AG067416.